OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of x2 = 0. a.no real root b.two distinct real roots c.one distinct real root d.three distinct real roots

OpenStudy (anonymous):

Whats the question...is it x^2=0 or x2=0...Plz type the qn correctly

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2=0

OpenStudy (anonymous):

Great... then the answer is C..There is only one root, x=0

OpenStudy (anonymous):

The reason for single root is D=0 here (i.e. b^2-4ac=0) in this case.

OpenStudy (anonymous):

thank you so much